How does a fixed cost per unit affect profit?
Answer #1

Answer

FIXED COST PER UNIT EFFECTS IN PROFIT

If the production increase fixed cost per unit decrease, it will leads to increase profit. If the production decrease fixed cost per unit increase, it will lead to decrease the profit.

For example

Fixed cost = $670

Selling price = $27

Variable cost = $5

1 COMPANY produced and SOLD 70 UNITS

Fixed cost per units = 670 / 70 = $9.57

Calculation of profit

Particulars $
Sales (70 × 27) 1,890
Less variable cost (70 × 5) 350
Contribution margin 1,540
Less fixed cost 670
Profit 870

2 IF PRODUCTION INCREASED TO 90 UNITS

Fixed cost per unit = 670 / 90 = $7.44

Calculation of profit

Particulars $
Sales (90 × 27) 2,430
Less variable cost (90 × 5) 450
Contribution margin 1,980
Less fixed cost 670
Profit 1,310

3 IF PRODUCTION IS DECREASED TO 55 UNITS

Fixed cost per unit = 670 / 55 = $12.18

Calculation of profit

Particulars $
Sales (55 × 27) 1,485
Less variable cost (55 × 5) 275
Contribution margin 1,210
Less fixed cost 670
Profit 540

The above calculations clearly indicate that the production increase from 70 unit to 90 units, it will lead to fixed cost per unit decrease from $9.57 to $7.44 and the profit increase from $870 to $1,310.

If the production decrease from 70 units to 55 units, it will lead to fixed cost per unit increase from $9.57 to $12.18 and the profit decrease from $870 to $540.

In short the change in fixed cost per unit will significantly effects the profit. If Fixed cost per unit increase profit decrease and fixed cost per unit decrease profit increase.

If YOU are like the answer please upvote.